# Optical pumping of alkali-metal vapor with hyperfine-resolved buffer gas pressure

### 06 Mar 2026 v1

@@ IV.3 Magnetic resonance enhancement
As described in Sec. III.3.3, the performance of atomic magnetometers can be reflected in the magnetic resonance linewidth $\Gamma_{2}$ and the maximum amplitude $Y_{\mathrm{max}}$. And the optical shift $\delta\Omega_{L}$ is also noticed. They are displayed in Fig. 8 to illustrate how different laser detunings enhance the magnetic resonance signal in the QHP regime.
In Fig. 8(a), $\Gamma_{2}(\Delta_{a}=0)$ and $\Gamma_{2}(\Delta_{b}=0)$ versus $\Phi$ are compared. More than well-known light-narrowing effect [2], we discover an extra narrowing effect induced by the resonant frequencies related to $\Delta_{F}$ for F=a,bF=a,\,b
$\Gamma_{2}(\Delta_{a}=0)>\Gamma_{2}(\Delta_{b}=0)$ (71)
generally both in the light-narrowing and light-broadening regimes. And the difference gradually disappears with larger $\Gamma_{\mathrm{brd}}$ in the HP limit, although the underlying mechanism remains operative. The narrowing effect can be exemplified by $\Gamma_{2}$ analytically derived in two polarization limits. In low polarization limit, the linewidth with first order correction is
$\Gamma^{(P\rightarrow 0)}_{2}$= $q_{\mathrm{ex}}(I)\Gamma_{\mathrm{ex}}+q_{\mathrm{SD}}(I)(\Gamma_{\mathrm{SD}}+R_{\mathrm{op}}\mathcal{Q}_{\mathrm{a}})$ (72)
where two slowing-down factors $q_{\mathrm{ex}}(I)$ and $q_{\mathrm{SD}}(I)$ are [22]
$q_{\mathrm{SD}}(I)=\frac{[I]^{2}}{2I^{2}+I+1},$ (73)
$q_{\mathrm{ex}}(I)=\frac{3[I]^{2}}{2I(2I-1)}.$ (74)
The traditional light-narrowing effect arises from higher order corrections including $1/(R_{\mathrm{op}}\mathcal{Q}_{\mathrm{a}})$ term. In high polarization limit, the linewidth is dominant by
$\Gamma^{(P\rightarrow 1)}_{2}$= $\frac{1}{[I]}R_{\mathrm{op}}\mathcal{Q}_{\mathrm{a}}+\frac{1}{2}\Gamma_{\mathrm{SD}}.$ (75)
In both Eqs. (72) and Eq. (75), the optical pumping contribution is determined by $R_{\mathrm{op}}Q_{a}={\Phi A_{0}\Gamma_{\mathrm{brd}}}/{(\Delta^{2}_{a}+\Gamma^{2}_{\mathrm{brd}})}$. Consequently, the cases of $\Delta_{a}=0$ and $\Delta_{b}=0$ result in additional broadening and narrowing, respectively. These behaviors originate from $|\tilde{\rho}_{1})$ dynamics governed by $\mathcal{G}_{\mathrm{WDA}}$, whose lowest eigenvalue in the $a$-subspace dominates $\Gamma_{2}$, as mentioned in Sec. III.3.
The narrowing effect induced by optical resonance with the $b$ multiplet significantly enhances the magnetic resonance signal, as is shown in Fig. 8(b) and Fig. 8(c). Evidently, compared with pumping at $\Delta_{a}=0$, the response near $\Delta_{b}=0$ is markedly superior, exhibiting a narrower resonance peak together with a substantially stronger amplitude. Such behaviors can be approximated and explained by the maximum amplitude of the out-of-phase quadrature $Y_{\mathrm{max}}\sim{P}/{\Gamma_{2}}$. For optical pumping in the QHP regime, there exists a regime of laser detuning in which increasing $P$ leads to stronger coherence source for $|\tilde{\rho}_{1})$, and meanwhile $\Gamma_{2}$ remains well protected or suppressed by the narrowing mechanism.
Additionally, the optical shift $\delta\Omega_{\mathrm{L}}$ of $\mathrm{S}$ described by Eq. (53) is included in Fig. 88. As an application, we emphasize that at the favored detuning $\Delta_{b}=0$
$\delta\Omega_{\mathrm{L}}(\Delta_{b}=0)=\frac{1}{[I]}\,\frac{\Phi A_{0}\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}}{\delta^{2}_{\mathrm{hf}}+\Gamma^{2}_{\mathrm{brd}}}.$ (76)
which means that $\delta\Omega_{\mathrm{L}}$ can be avoided or minimized [5], at specific laser detunings without sacrifice of the magnetic resonance linewidth and amplitude.
@@ V Conclusions
We developed a theoretical framework for optical pumping of alkali-metal atoms in buffer-gas cells in the QHP regime, where the collisional broadening $\Gamma_{\mathrm{brd}}$ is comparable to the GS hyperfine splitting, and far exceeds any other splitting or broadening. Via the master equation in Liouville space including various optical and collisional mechanisms, we studied optical pumping with the $F=a$ and $F=b$ hyperfine multiplet resolved and establish a compact relation linking QHP superoperators to their HP counterparts, providing a unified theoretical framework. Moreover, we investigate the steady state beyond STD and point out the validity of the STD limit under typical conditions. To apply these treatments to atomic megnetometers, We further analyzed observables including the electron spin polarization, the absorption cross section and the magnetic resonance linewidth. Finally, We find and explain that in the QHP regime the magnetic resonance is significantly enhanced, exhibiting both linewidth narrowing and increased resonance amplitude at laser detuning near $\Delta_{b}=0$.
In summary, our study provides a deep understanding of optical pumping in the QHP regime and offers useful theoretical tools with practical applicability to atomic sensors (e.g., atomic magnetometers), operating under realistic buffer-gas pressures. It also establishes a foundation for extending the present framework to incorporate additional physical effects, such as atomic diffusion.
VI Acknowledgments
@@
Appendix A BLOCKS IN lIOUVILLE SPACE
@@
In Liouville-space calculations, an appropriate ordering of the basis states is essential for exposing the block structure of large superoperators. Such a decomposition not only improves computational efficiency but also highlights the underlying physical organization of the dynamics. In this Appendix, we introduce a block scheme constructed by grouping density-matrix elements according to their characteristic evolution frequencies [22], with emphasis on the blocks associated with the Zeeman subspaces.
The projection operator of the GS $k$th-order Zeeman coherence $\mathcal{P}^{\{k\}}_{\mathrm{Z}}$ is introduced as

### 14 Apr 2026 v2

@@ IV.3 Magnetic resonance enhancement
For polarized atoms under RF driving, we specifically investigate the impact of QHP-specific optical pumping on $\Gamma_{2}$, which governs the performance of the intrinsic magnetic resonance discussed in Sec. III.3.3. This impact is explicitly described by the light-induced linewidth variation
$\delta\Gamma_{2}=\Gamma_{2}-\Gamma_{2,\,\mathrm{col}}.$ (59)
Here, $\Gamma_{2,\,\mathrm{col}}$ serves as the baseline linewidth obtained at vanishing photon flux i.e, $P\rightarrow 0$. For nuclear spin $I$, it takes the form determined by the spin-destruction and spin-exchange collision rates
$\Gamma_{2,\,\mathrm{col}}=$ $q_{\mathrm{SD}}(I)\Gamma_{\mathrm{SD}}+q_{\mathrm{ex}}(I)\Gamma_{\mathrm{ex}},$ (60)
with $q_{\mathrm{SD}}(I)$ and $q_{\mathrm{ex}}(I)$ the slowing-down factors [22]
$q_{\mathrm{SD}}(I)=\frac{(2I+1)^{2}}{2I^{2}+I+1},$ (61)
$q_{\mathrm{ex}}(I)=\frac{3(2I+1)^{2}}{2I(2I-1)}.$ (62)
The light-independent form of $\Gamma_{2,\,\mathrm{col}}$ remains invariant to buffer gas pressure, which only linearly scales $\Gamma_{\mathrm{SD}}$.
For representative QHP and HP conditions, Figure 8(a) plots $\delta\Gamma_{2}$ at optical resonance against the photon flux $\Phi$. For the traditional HP case (dashed line), the $F$-unresolved behavior of $\delta\Gamma_{2}$ is expected. Under rapid spin-exchange conditions, it initially decreases and then increases, reproducing the classical linewidth narrowing and broadening phenomena in specific polarization ranges [22, 2]. While $\delta\Gamma_{2}$ in the QHP regime (solid lines) preserves this behavior with increasing $\Phi$, it manifests pronounced $F$-resolved features. Across the predominant range of photon fluxes, the linewidth at $\Delta_{b}=0$ (blue line) is consistently narrower than that at $\Delta_{a}=0$ (red line), indicating that pumping at the $F=b$ resonance possesses an extra linewidth suppression capability, which fundamentally enables the magnetic resonance to be enhanced.
We explain this narrowing mechanism through the asymmetric absorption of $F=a$ and $F=b$ and its impact on the superoperator $\mathcal{G}_{\mathrm{WDA}}$ governing the relaxation of $|\tilde{\rho}_{1})$. Physically, for polarized atoms, the $F=a$ multiplet provides the dominant population imbalance that drives the coherence $|\tilde{\rho}_{1})$. Therefore, the overall relaxation of $|\tilde{\rho}_{1})$, which determines the resonance linewidth, is fundamentally dictated by the specific $\mathcal{G}_{\mathrm{WDA}}$ block corresponding to the $F=a$ level, i.e., $\mathcal{P}_{a}^{\{1\}}\mathcal{G}_{\mathrm{WDA}}\mathcal{P}_{a}^{\{1\}}$. Consistently, the optical pumping contribution $\mathcal{A}_{\mathrm{op}}]_{1,1}$ to this block originates from absorption of the $F=a$ level. This drives the conclusion that the direct light-induced relaxation incorporated into the linewidth is scaled by the weight factor $\mathcal{Q}_{a}$. When the pumping light is tuned near $\Delta_{b}=0$, $\mathcal{Q}_{a}$ attains values much smaller than 1 (see Fig. 2), manifesting as a practical suppression of decoherence during the optical pumping process. Conversely, the $F=a$ optical resonance exhibits $\mathcal{Q}_{a}>1$, resulting in a relative broadening of the linewidth. The $\mathcal{Q}_{a}$-scaling effect becomes particularly evident under intense light ($P\to 1$), where the pumping process dominates the linewidth behavior. In this case, the linewidth asymptotically approaches
$\Gamma_{2}=$ $\frac{1}{2I+1}A\Phi\mathcal{L}(\Delta_{a})+\frac{1}{2}\Gamma_{\mathrm{SD}},$ (63)
obtained by replacing the $F$-unresolved pumping rate in the HP counterpart [22] by $R_{\mathrm{op}}Q_{a}=A\Phi\mathcal{L}(\Delta_{a})$. This illustrates that the $F=b$ optical resonance essentially retards the growth of linewidth with increasing $\Phi$.
Sharing the same physical origin as the linewidth, the light modifies the Larmor frequency by the optical shift $\delta\omega_{\mathrm{L}}$ expressed as
$\delta\omega_{\mathrm{L}}$ $=\frac{\Delta_{a}}{\Gamma_{\mathrm{brd}}}\frac{A\Phi\mathcal{L}(\Delta_{a})}{2I+1}.$ (64)
While crossing zero at $\Delta_{a}=0$, detuning away from this point results in a residual shift. Fortunately, under favorable $\Gamma_{\mathrm{brd}}$ conditions, the optical shift can be attenuated by operating near the $F=b$ resonance where the perturbation to Larmor precession within the $F=a$ level is negligible. Notably, these QHP-specific theoretical predictions regarding the linewidth $\Gamma_{2}$ and light shift $\delta\omega_{\mathrm{L}}$ accord with previous experimental observations [5].
Furthermore, these features in physics are directly reflected in the observable $X$ and $Y$ signals, especially at large $\Phi$ as shown in Figs. 8(b) and 8(c). Although operating at $\Delta_{b}=0$ introduces a modest optical shift, it serves as a favorable trade-off for the overall magnetic resonance performance. When pumping at $\Delta_{b}=0$ (blue lines), the magnetic resonance spectra feature an enhanced signal amplitude compared to the $\Delta_{a}=0$ condition (red lines). We understand this enhancement behavior using the maximum resonance amplitude of the quadrature $Y$ as a representative metric. Due to its Lorentzian line shape, the resonance peak is evaluated by $Y_{\mathrm{max}}\sim P/\Gamma_{2}$, where $P$ acts as the fundamental source of $|\tilde{\rho}_{1})$. Coupling the linewidth suppression discussed above with the behavior of $P$ detailed in Sec. IV.2, we proceed to analyze $Y_{\mathrm{max}}$ across $\Phi$ domains. When $\Phi<\Phi^{\prime}_{\mathrm{eq}}$, the competing effects of $P$ and $\Gamma_{2}$ reach a close draw for $Y_{\mathrm{max}}$ between pumping at $\Delta_{a}=0$ and $\Delta_{b}=0$. In contrast, when $\Phi>\Phi^{\prime}_{\mathrm{eq}}$, the $F=b$ resonance not only maintains a higher $P$ to act as a powerful coherence source (enlarging the numerator), and meanwhile, its $\mathcal{Q}_{a}$-suppression effect protects $\Gamma_{2}$ from severe light broadening (lowering the denominator), thereby yielding a substantially enhanced magnetic resonance performance.
It should be noted that the physical demonstrations above primarily contrast the resonant pumping conditions where the $F$-resolved absorption and optical pumping behavior are most pronounced. In practical atomic magnetometer applications, while off-resonant pumping is still worth exploring for globally optimizing sensitivity, the $F=b$ near-resonance pumping configuration remains a preferred pathway due to the underlying physical mechanisms discussed above. In this section, the conclusions rooted in principles can be extended beyond the displayed $^{87}\text{Rb}$ system to other alkali-metal species.
@@ V Conclusions
We developed a theoretical framework for optical pumping of alkali-metal atoms in buffer-gas cells operating in the QHP regime, where the collisional broadening $\Gamma_{\mathrm{brd}}$ is comparable to the GS hyperfine splitting $\omega^{\{\mathrm{g}\}}_{\mathrm{hf}}$. By incorporating the weighting factor $Q_{F}$ to account for asymmetric absorption, we derived the QHP-specific optical pumping superoperator within the Liouville-space master equation framework. Its $F$-resolved form established a compact relation linking to the HP counterpart, and facilitates both physical analysis and numerical evaluation in the QHP regime.
Using this unified approach, we demonstrated that QHP dynamics are governed by the combined effects of population redistribution across both $F$ multiplets and $m_{F}$ levels. We presented and discussed these $F$-pumping and $m_{F}$-pumping effects under different light polarizations. Furthermore, we examined the resulting steady-state populations that deviate from the standard STD and pointed out the validity of the STD approximation under these conditions.
Regarding observables with QHP corrections, we systematically evaluated three interconnected quantities under $F$-resolved pumping: the absorption cross section $\sigma$, the electron spin polarization $P$, and the intrinsic magnetic resonance linewidth $\Gamma_{2}$. The characteristic behaviors of absorption and polarization are highlighted by features such as the convergent integral area under linear polarization, alongside the distinct equal-absorption and equal-polarization points under circular polarization. Applying these insights to atomic magnetometry, we identified an optimal operating scheme, demonstrating that pumping near the $F=b$ resonance comprehensively enhances performance by suppressing the linewidth and amplifying the signal amplitude.
In summary, our study provides a profound understanding of the QHP-specific optical pumping and offers useful theoretical tools with practical applicability to atomic sensors (e.g. atomic magnetometers) operating under realistic buffer-gas pressures. The methodology developed herein also establishes a foundation for integrating further physical effects in the QHP regime, including atomic diffusion.
ACKNOWLEDGMENTS
@@
DATA AVAILABILITY
The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
Appendix A THE QHP CORRECTION IN LIOUVILLE SPACE
With the QHP correction Eq. (19), we offer detailed proof for the $F$-resolved form of superoperators $\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}$ Eq. (22) and $\mathcal{A}^{\{\mathrm{e}g\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}$ Eq. (27), and also show that this compact form are not available for hyperfine coherences between $F=a$ and $F=b$ multiplet.
To simplify the depopulation superoperator, we utilize the identities in Liouville space [9],
${X}^{\textcopyright}={X}^{\flat}-{X^{\dagger\sharp}},$ (65)
$(AB)^{\flat}=A^{\flat}B^{\flat},$ (66)
$(AB)^{\sharp}=B^{\sharp}A^{\sharp},$ (67)
with operators A,BA,\,B and $X$ in Hilbert space. Therefore, arising from the $F$-resolved $\delta H^{\{\mathrm{g}\}}$ in Eqs. (20) &\& (21), $\mathcal{A}_{\mathrm{dp}}$ Eq. (11) becomes
$\mathcal{A}_{\mathrm{dp}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}\Pi_{F}^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}\Pi_{F}^{\sharp}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right).$ (68)
The relation between projection operators in Hilbert and Liouville spaces is
$\Pi_{F}^{\flat}$ $=\mathcal{P}_{F}+\mathcal{P}_{F\to F^{\prime}},$ (69)
$\Pi_{F}^{\sharp}$ $=\mathcal{P}_{F}+\mathcal{P}_{F^{\prime}\to F}.$ (70)
Here, we define the projection operator of hyperfine coherence as ($F\neq F^{\prime}$)
$\mathcal{P}_{F\to F}$ $=\sum_{m_{F},m^{\prime}_{F}}\left|F,m_{F};F^{\prime},m^{\prime}_{F}\right)\left(F,m_{F};F^{\prime},m^{\prime}_{F}\right|,$ (71)
$\mathcal{P}_{\mathrm{hf}}$ $=\mathcal{P}_{a\to b}+\mathcal{P}_{b\to a}.$ (72)
These properties of projection operators enable us to derive
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right)\mathcal{P}_{F},$ (73)
$\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{hf}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{\left(V^{\dagger}V\right)^{\flat}\mathcal{P}_{F\to F^{\prime}}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{\left(V^{\dagger}V\right)^{\sharp}\mathcal{P}_{F^{\prime}\to F}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right).$ (74)
Here, the first line proves that, after the QHP correction, $\mathcal{A}_{\mathrm{dp}}\mathcal{P}_{\mathrm{Z}}$ acquires an $F$-resolved structure, allowing it to be decomposed into separate contributions from the $F=a$ and $F=b$ levels, in close analogy with $\delta H^{\{\mathrm{g}\}}$. In contrast, the second line indicates that the contributions from positive- and negative-frequency hyperfine coherences to the depopulation dynamics cannot be cast into a unified form in the same way. For further simplification, the weight factor $Q_{F}$ Eq. (23) is introduced
$Q_{F}\equiv\frac{\pi|\Omega_{\mathrm{I}}|^{2}\mathcal{L}(\Delta_{F})}{R_{\mathrm{op}}}=\frac{\mathcal{L}(\Delta_{F})}{x_{a}\mathcal{L}(\Delta_{a})+x_{b}\mathcal{L}(\Delta_{b})}.$ (75)
Therefore, we can employ the symmetry property of the dipole operator $\hat{\bm{e}}^{*}\cdot\bm{d}\bm{d}^{\dagger}\cdot\hat{\bm{e}}=1/2-K\mathbf{s}\cdot\mathbf{S}$ and obtain the compact form Eq. (22) expressed in terms of spin operators and projection operators.
We now turn to the optical excitation superoperator $\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}$. Following the same procedure as above, we focus on its $F$-resolved form Eq. (27). Starting from the definition Eq. (14), we obtain
$\mathcal{A}^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}=$ $\frac{i}{R_{\mathrm{op}}}\sum_{F}\left(\frac{V^{\dagger\sharp}V^{\flat}}{\Delta_{F}+i\Gamma_{\mathrm{brd}}}-\frac{V^{\dagger\sharp}V^{\flat}}{\Delta_{F}-i\Gamma_{\mathrm{brd}}}\right)\mathcal{P}_{F}.$ (76)
As in the similar case of Eq. (74), this form also does not extend to the hyperfine coherence. Here, we utilize the property ${B}^{\mathrm{T}}\otimes A=A^{\flat}B^{\sharp}$ [9]. With the factor $Q_{F}$, $A^{\{\mathrm{eg}\}}_{\mathrm{p}}\mathcal{P}_{\mathrm{Z}}$ is expressed as Eq. (27). In the HP limit $Q_{F}\to 1$, the excitation dynamics becomes $F$-unresolved, and the superoperator reduces to
[…]
In summary, the QHP corrections derived for the various superoperators consistently follow the same $Q_{F}$-weighted and $F$-resolved formalism, thereby establishing a clear connection to the traditional HP limit. This structural equivalence is essential for constructing a compact expression of the optical pumping superoperator $\mathcal{A}_{\mathrm{op}}$.
Appendix B FREQUENCY BLOCKS IN ZEEMAN SUBSPACE
@@
In Liouville-space calculations for the master equation, an appropriate ordering of the basis states is essential for exposing the block structure of large superoperators. Such a structure not only improves computational efficiency but also highlights the underlying physical organization of the dynamics. A block group scheme is constructed in Zeeman subspace, i.e., $\mathrm{span}\{|Fm_{F},Fm^{\prime}_{F})\}$ according to the characteristic evolution frequencies of the density-matrix elements [22].
The projection operators of the GS $k$th-order Zeeman coherence are introduced as